OpenStudy (ujjwal):

p and q are odd numbers, prove that the quadratic equation \(x^2+px+q=0\) has 1) no integer solution 2) no rational solution

OpenStudy (amistre64):

use the quadratic formula ....

OpenStudy (amistre64):

\[x=\frac{-(2k+1)\pm\sqrt{(2k+1)^2-4(2k+1)(2n+1)}}{2(2n+1)}\]

OpenStudy (amistre64):

i misplaced my p and q ...

OpenStudy (amistre64):

\[x=\frac{-(2k+1)\pm\sqrt{(2k+1)^2-4(1)(2n+1)}}{2}\]

OpenStudy (ujjwal):

i guess, basically we need to prove that the discriminant is never a positive perfect square number. Don't know if that's a correct approach..

OpenStudy (amistre64):

yes, and recall that k,n in Z

OpenStudy (amistre64):

if:\[4k^2+4k-8n-8+1\ge0\] then there is a chance for it, so as long as we can show the other stuff notice that is an odd number to start with ..... not that there arent odd perfect squares, but it might reduce the work load

OpenStudy (ujjwal):

yep. and it's \(4k^2+4k-8n-4+1\geq 0\) and 9 is an odd number and a perfect square. are there no other odd perfect squares?

OpenStudy (amistre64):

15, 49, 81, odd^2 = odd 4(2n-1) >= 4k^2 + 4k + 1 2n-1 >= k^2 + k + 1/4

OpenStudy (ujjwal):

squares of all odd numbers are odd

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

2n+1 >= k^2 + k + 1/4 2n + 1 >= (k +1/2)^2 n >= [(k +1/2)^2 - 1] /2

OpenStudy (ujjwal):

if that's the case, the discriminant is positive. but we don't even know if that's the case or not or any such thing..

OpenStudy (ujjwal):

I don't get it much..

OpenStudy (amistre64):

do we have to try a direct approach? or can we use contradiction or contrapositive?

OpenStudy (ujjwal):

I don't have much idea.. Any which way, this has to be proved..

OpenStudy (anonymous):

try to show that the discriminant of quadratic can not be a perfect square

OpenStudy (ujjwal):

exactly.. but i don't know how?

OpenStudy (amistre64):

if sqrt(p^2-4q) is odd, then we have an integer if sqrt(p^2-4q) is a perfect square, then we have a rational

ganeshie8 (ganeshie8):

see if it works :- say, \(\large x^2+px+q = (x-\alpha)(x-\beta)\) comparing coeffecients both sides, \(\large \alpha + \beta = -p\) ----(1) \(\large \alpha \beta = q\) ----- (2) substitute \(\beta = -p- \alpha\) from (1) in (2) \(\large -\alpha(p + \alpha) = q\) right hand side is a odd, but left hand side is even. contradiction - so integer roots are not possible

OpenStudy (amistre64):

a+b is odd if a and b are odd odd(odd+odd) = odd ? odd(even) = odd ? 2n(2k+1) = 2(2nk+n) is even

OpenStudy (ujjwal):

i get it @ganeshie8 . Thanks! But then what about the irrational thing?

OpenStudy (amistre64):

3+1 = 4 3 + 3 = 6 odd + odd = even

OpenStudy (ujjwal):

i mean the second question which is asking to prove that the solution is not rational.

OpenStudy (amistre64):

im outta time for today ... good luck

ganeshie8 (ganeshie8):

oh, second q looks tricky... @mukushla

OpenStudy (ujjwal):

\(\alpha +\beta =p\) which is odd, so, among \(\alpha\) and \(\beta\) one is odd while the other is even. but then, \(\alpha\beta=q\) left side is even, right side is odd contradictory

ganeshie8 (ganeshie8):

u made it look much simpler :)

ganeshie8 (ganeshie8):

### amistre lines above : if sqrt(p^2-4q) is odd, then we have an integer if sqrt(p^2-4q) is a perfect square, then we have a rational should work ?

OpenStudy (anonymous):

oh guys, i was out...let\[\Delta=p^2-4q=n^2\]\(n\) is odd, let \(p=2k+1 , \ n=2m+1\)\[(p-n)(p+n)=4q\]\[(k-m)(k+m+1)=q\]left hand side is even and right hand side is odd hence both parts of question proved

OpenStudy (ujjwal):

Thanks a lot! You made it a lot simpler!

OpenStudy (anonymous):

very welcome bro :)